Q.) A bomb of mass 300kg fragmented into two parts :one having mass 200kg moving with a velocity of 16m/s² and other having mass 100kg.Find out total kinetic energy produced in this bombarding


1
Expert's answer
2020-04-13T10:15:02-0400

The speed of second part of a bomb we can fin using the law of conservation of momentum

"0=m_1v_1-m_2v_2."

Hence,

"v_2=m_1v_1/m_2=200\times 16/100=32\:\rm m/s."

The total kinetic energy

"K=\frac{m_1v_1^2}{2}+\frac{m_2v_2^2}{2}=\\
=\frac{200\times 16^2}{2}+\frac{100\times 32^2}{2}=76800\:\rm J."